# Denoising Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

### v3

@@ Abstract
We propose denoising diffusion variational inference (DDVI), a black-box variational inference algorithm for latent variable models which relies on diffusion models as flexible approximate posteriors. Specifically, our method introduces an expressive class of diffusion-based variational posteriors that perform iterative refinement in latent space; we train these posteriors with a novel regularized evidence lower bound (ELBO) on the marginal likelihood inspired by the wake-sleep algorithm. Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks. We find that DDVI improves inference and learning in deep latent variable models across common benchmarks as well as on a motivating task in biology—inferring latent ancestry from human genomes—where it outperforms strong baselines on the Thousand Genomes dataset.
@@ 1 Introduction
$\log\;p_{\bm{\theta}}(\mathbf{x})\geq\max_{\bm{\phi}}\;\;\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log\;p_{\bm{\theta}}(\mathbf{x},\mathbf{z})-\log\;q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})\right]:=\max_{\bm{\phi}}\;\;\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}}),$ (1)
in which we approximate the marginal likelihood $\log\;p_{\bm{\theta}}(\mathbf{x})$ of a latent variable model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ with an evidence lower bound $\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ that is a function of an approximate variational posterior $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$. We assume that $p_{\bm{\theta}}$ factorizes as $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\bm{\theta}}$ include Bayesian networks, topic models [4], variational autoencoders (VAEs), and broad classes of $p_{\bm{\theta}}$ defined via modern probabilistic programming frameworks [9].
Maximizing $\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ over ${\bm{\phi}}$ yields a variational posterior $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ that approximates $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ as well as a tight bound on $\log\;p_{\bm{\theta}}(\mathbf{x})$ that serves as a learning objective for $p_{\bm{\theta}}$. The approximation gap $\log\;p_{\bm{\theta}}(\mathbf{x})-\max_{\bm{\phi}}\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ equals precisely $\min_{\bm{\phi}}\text{KL}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$, which motivates the design of expressive classes of posteriors $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ that reduce this gap. Recent efforts leverage modern generative models—including normalizing flows [28, 16] and generative adversarial networks [8, 20]—as expressive model families for $q_{\bm{\phi}}$ that tighten the ELBO.
This work seeks to further improve variational inference via expressive posteriors based on diffusion models [12, 37]. Diffusion methods have become the de-facto standard for high-quality image synthesis [30]. Here, we use diffusion in latent space to parameterize $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$. We train this distribution with a denoising diffusion-like objective that does not involve adversarial training [20] or constrained invertible normalizing flow architectures [16]. Samples from $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ are obtained via iterative refinement of $\mathbf{z}$, starting from a Gaussian distribution, and gradually forming one that is multi-modal and complex.
@@
Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm defined by a class of approximate posterior distribution based on diffusion as well as a learning objective inspired by the wake-sleep algorithm [11] that can be interpreted as a regularized variational inference. We also derive extensions of our method to semi-supervised learning and clustering.
Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks. We evaluate DDVI on synthetic benchmarks and on a real problem in biological data analysis—inferring human ancestry from genetic data. Our method outperforms strong baselines on the Thousand Genomes dataset [35] and learns a low-dimensional latent space that preserves biologically meaningful structure [10].
@@ 2 Background
Latent variable models (LVMs) $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ are usually fit by optimizing the evidence lower bound (ELBO)
$\log p_{\bm{\theta}}(\mathbf{x})\geq\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z})),$ (2)
which serves as a tractable surrogate for the marginal log-likelihood (MLL). The gap between the MLL and the ELBO equals precisely $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO [14].
Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods [19], MCMC-based methods [31], normalizing flows [28]—or improved learning objectives—e.g., adversarial or sample-based losses [20, 48, 33, 34].
The wake-sleep algorithm [11] optimizes an alternative objective
$\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})),$ (3)
in which the KL divergence term is reversed. The learning procedure for wake-sleep involves alternating between "wake" phases where the recognition model is updated and "sleep" phases where the generative model is refined.
@@
A diffusion model is defined via a user-specified noising process $q$ that maps data $\mathbf{x}_{0}$ into a sequence of $T$ variables $\mathbf{y}_{1:T}=\mathbf{y}_{1},...,\mathbf{y}_{T}$ that represent increasing levels of corruption to $\mathbf{x}_{0}$. We obtain $\mathbf{y}_{1:T}$ by applying a Markov chain $q(\mathbf{y}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where we define $\mathbf{y}_{0}=\mathbf{x}_{0}$ for convenience. When $\mathbf{x}_{0}$ is a continuous vector, a standard choice of transition kernel is $q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{{\bm{\alpha}}_{t}}\mathbf{y}_{t-1},\sqrt{1-{\bm{\alpha}}_{t}}\mathbf{I})$, which is a Gaussian centered around a copy of $\mathbf{y}_{t-1}$ to which we added noise following a schedule $10<{\bm{\alpha}}_{1}<{\bm{\alpha}}_{2}<...<{\bm{\alpha}}_{T}=$ 1.
@@
A diffusion model can then be represented as a latent variable distribution $p(\mathbf{x}_{0},\mathbf{y}_{1:T})$ that factorizes as $p(\mathbf{x}_{0},\mathbf{y}_{1:T})=p(\mathbf{y}_{T})\prod_{t=0}^{T-1}p_{\bm{\theta}}(\mathbf{y}_{t}\mid\mathbf{y}_{t+1})$ (again using $\mathbf{y}_{0}$ as shorthand for $\mathbf{x}_{0}$). This model seeks to approximate the reverse of the forward diffusion $q$ and map noise $\mathbf{y}_{T}$ into data $\mathbf{x}_{0}$.
@@
The true reverse of the process $q$ cannot be expressed in closed form; as such, we parameterize $p_{\bm{\theta}}$ with ${\bm{\theta}}$ trained by maximizing the ELBO:
$\log p_{\bm{\theta}}(\mathbf{x}_{0})\geq\mathbb{E}_{q}\Bigr{[}\log p_{\bm{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{1})-\sum_{t=2}^{T}D_{\mathrm{KL}}(q_{t}||p_{t})\Bigr{]}-D_{\mathrm{KL}}(q(\mathbf{x}_{T}|\mathbf{x}_{0})||p(\mathbf{x}_{T}))$ (4)
where $q_{t},p_{t}$ denote the distributions $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ and $p_{\bm{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, respectively.
@@
3 Variational Inference With Denoising Diffusion Models
We introduce denoising diffusion variational inference (DDVI), which improves variational inference with diffusion-based techniques. The goal of DDVI is to fit a latent variable model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$. We assume that $p_{\bm{\theta}}$ factorizes as $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\bm{\theta}}$ include Bayesian networks and variational autoencoders (VAEs) [14]. Our approach is comprised of two key components: a family of approximate posteriors $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ based on diffusion and a learning objective that can be seen as a regularized ELBO. The $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ iteratively refines latents $\mathbf{z}$, starting from a Gaussian distribution. The learning objective trains $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to reverse a used-specified forward diffusion process.
3.1 Diffusion-Based Variational Posteriors
@@
We choose the latent $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ to be a vector of $T$ variables that represent progressively simplified versions of $\mathbf{z}$, with $\mathbf{y}_{T}$ corresponding to a simple distribution (e.g., a Gaussian). The model $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\bm{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$ transforms $\mathbf{y}_{T}$ into $\mathbf{z}$ via iterative refinement. To sample from $q_{\bm{\phi}}$, we first sample $\mathbf{y}_{T}$—this is an easier task since we can define $\mathbf{y}_{T}$ to have a simple (e.g., Gaussian) distribution—and then by sampling from the denoising model $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\bm{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$.
@@
We define the relationship between $\mathbf{y}$ and $\mathbf{z}$ via a forward diffusion process $r(\mathbf{y}|\mathbf{z})=r(\mathbf{y}_{1:T}|\mathbf{z})=r(\mathbf{y}_{1}|\mathbf{z})\prod_{t=1}^{T-1}r(\mathbf{y}_{t+1}|\mathbf{y}_{t})$, which transforms $\mathbf{z}$—the latent whose intractable posterior we seek to approximate—into $\mathbf{y}_{T}$, whose posterior is easier to model. Examples of $r$ include Gaussian forward diffusion processes and discrete noising processes [2]. The model $q_{\bm{\phi}}$ is trained to approximately reverse this forward diffusion process.
3.2 The Wake-Sleep Regularized ELBO

### v4

@@ Abstract
We propose denoising diffusion variational inference (DDVI), a black-box variational inference algorithm for latent variable models which relies on diffusion models as flexible approximate posteriors. Specifically, our method introduces an expressive class of diffusion-based variational posteriors that perform iterative refinement in latent space; we train these posteriors with a novel regularized evidence lower bound (ELBO) on the marginal likelihood inspired by the wake-sleep algorithm. Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks. We find that DDVI improves inference and learning in deep latent variable models across common benchmarks as well as on a motivating task in biology—inferring latent ancestry from human genomes—where it outperforms strong baselines on 1000 Genomes dataset.
@@ 1 Introduction
$\log\;p_{\boldsymbol{\theta}}(\mathbf{x})\geq$ $\max_{\boldsymbol{\phi}}\;\;\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log\;p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})-\log\;q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})\right]$
$:=\max_{\boldsymbol{\phi}}\;\;\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}}),$ (1)
in which we approximate the marginal likelihood $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})$ of a latent variable model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ with an evidence lower bound $\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ that is a function of an approximate variational posterior $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$. We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks, topic models (Blei, Ng, and Jordan 2003), variational autoencoders (VAEs), and broad classes of $p_{\boldsymbol{\theta}}$ defined via modern probabilistic programming frameworks (Gordon et al. 2014).
Maximizing $\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ over ${\boldsymbol{\phi}}$ yields a variational posterior $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ that approximates $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$ as well as a tight bound on $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})$ that serves as a learning objective for $p_{\boldsymbol{\theta}}$. The approximation gap $\log\;p_{\boldsymbol{\theta}}(\mathbf{x})-\max_{\boldsymbol{\phi}}\text{ELBO}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}})$ equals precisely $\min_{\boldsymbol{\phi}}\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$, which motivates the design of expressive classes of posteriors $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ that reduce this gap. Recent efforts leverage modern generative models—including normalizing flows (Rezende and Mohamed 2015; Kingma et al. 2016) and generative adversarial networks (Goodfellow et al. 2014; Makhzani et al. 2015)—as expressive model families for $q_{\boldsymbol{\phi}}$ that tighten the ELBO.
This work seeks to further improve variational inference via expressive posteriors based on diffusion models (Ho, Jain, and Abbeel 2020; Song et al. 2020). Diffusion methods have become the de-facto standard for high-quality image synthesis (Rombach et al. 2022; Gokaslan et al. 2024). Here, we use diffusion in latent space to parameterize $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$. We train this distribution with a denoising diffusion-like objective that does not involve adversarial training (Makhzani et al. 2015) or constrained invertible normalizing flow architectures (Kingma et al. 2016). Samples from $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ are obtained via iterative refinement of $\mathbf{z}$, starting from a Gaussian distribution, and gradually forming one that is multi-modal and complex.
@@
Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm defined by a class of approximate posterior distribution based on diffusion and a learning objective inspired by the wake-sleep algorithm (Hinton et al. 1995) that implements regularized variational inference. We also derive extensions of our method to semi-supervised learning and clustering.
Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks.
We evaluate DDVI on synthetic benchmarks and on a real problem in biological data analysis—inferring human ancestry from genetic data. Our method outperforms strong baselines on 1000 Genomes dataset (Siva 2008) and learns a low-dimensional latent space that preserves biologically meaningful structure (Haghverdi, Buettner, and Theis 2015).
@@ 2 Background
Latent variable models (LVMs) $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$ are usually fit by optimizing the evidence lower bound (ELBO)
$\log p_{\boldsymbol{\theta}}(\mathbf{x})\geq\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z})),$
which serves as a tractable surrogate for the marginal log-likelihood (MLL). The gap between the MLL and the ELBO equals precisely $\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO (Kingma and Welling 2013).
Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods (Maaløe et al. 2016), MCMC-based methods (Salimans, Kingma, and Welling 2015), normalizing flows (Rezende and Mohamed 2015)—or improved learning objectives—e.g., adversarial or sample-based losses (Makhzani et al. 2015; Zhao, Song, and Ermon 2017; Si, Bishop, and Kuleshov 2022; Si et al. 2023).
The wake-sleep algorithm (Hinton et al. 1995) optimizes an alternative objective
$\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]-\text{KL}(p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})),$
in which the KL divergence term is reversed. The learning procedure for wake-sleep involves alternating between ”wake” phases where the recognition model is updated and ”sleep” phases where the generative model is refined.
@@
A diffusion model is defined via a user-specified noising process $q$ that maps data $\mathbf{x}_{0}$ into a sequence of $T$ variables $\mathbf{y}_{1:T}=\mathbf{y}_{1},...,\mathbf{y}_{T}$ that represent increasing levels of corruption to $\mathbf{x}_{0}$. We obtain $\mathbf{y}_{1:T}$ by applying a Markov chain $q(\mathbf{y}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where we define $\mathbf{y}_{0}=\mathbf{x}_{0}$ for convenience. When $\mathbf{x}_{0}$ is a continuous vector, a standard choice of transition kernel is $q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{{\boldsymbol{\alpha}}_{t}}\mathbf{y}_{t-1},\sqrt{1-{\boldsymbol{\alpha}}_{t}}\mathbf{I})$, which is a Gaussian centered around a copy of $\mathbf{y}_{t-1}$ to which we added noise following a schedule $10<{\boldsymbol{\alpha}}_{1}<{\boldsymbol{\alpha}}_{2}<...<{\boldsymbol{\alpha}}_{T}=$ 1.
@@
A diffusion model can then be represented as a latent variable distribution $p(\mathbf{x}_{0},\mathbf{y}_{1:T})$ that factorizes as $p(\mathbf{x}_{0},\mathbf{y}_{1:T})=p(\mathbf{y}_{T})\prod_{t=0}^{T-1}p_{\boldsymbol{\theta}}(\mathbf{y}_{t}\mid\mathbf{y}_{t+1})$ (again using $\mathbf{y}_{0}$ as shorthand for $\mathbf{x}_{0}$). This model seeks to approximate the reverse of the forward diffusion $q$ and map noise $\mathbf{y}_{T}$ into data $\mathbf{x}_{0}$.
@@
The true reverse of the process $q$ cannot be expressed in closed form; as such, we parameterize $p_{\boldsymbol{\theta}}$ with ${\boldsymbol{\theta}}$ trained by maximizing the ELBO:
$\log p_{\boldsymbol{\theta}}(\mathbf{x}_{0})\geq$ $\mathbb{E}_{q}\Bigr{[}\log p_{\boldsymbol{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{1})-\sum_{t=2}^{T}\text{KL}(q_{t}||p_{t})\Bigr{]}$ (2)
$-\text{KL}(q(\mathbf{x}_{T}|\mathbf{x}_{0})||p(\mathbf{x}_{T}))$
where $q_{t},p_{t}$ denote the distributions $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ and $p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, respectively.
3 Variational Inference With
Denoising Diffusion Models
@@ 3 Variational Inference With
We introduce denoising diffusion variational inference (DDVI), which improves variational inference with diffusion-based techniques.
The goal of DDVI is to fit a latent variable model $p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})$. We assume that $p_{\boldsymbol{\theta}}$ factorizes as $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})p_{\boldsymbol{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\boldsymbol{\theta}}$ include Bayesian networks and variational autoencoders (VAEs) (Kingma and Welling 2013).
Our approach is comprised of three components:
  1. 1.
    A modeling family of approximate posteriors $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ based on diffusion;
  2. 2.
    A learning objective formed by a regularized ELBO;
  3. 3.
    An optimization algorithm inspired by wake-sleep.
The $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ iteratively refines latents $\mathbf{z}$, starting from a Gaussian distribution. The learning objective trains $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ to reverse a used-specified forward diffusion process.
3.1 Modeling Family: Diffusion-Based Posteriors
@@
We choose the latent $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ to be a vector of $T$ variables that represent progressively simplified versions of $\mathbf{z}$, with $\mathbf{y}_{T}$ corresponding to a simple distribution (e.g., a Gaussian). The model $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\boldsymbol{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$ transforms $\mathbf{y}_{T}$ into $\mathbf{z}$ via iterative refinement. To sample from $q_{\boldsymbol{\phi}}$, we first sample $\mathbf{y}_{T}$—this is an easier task since we can define $\mathbf{y}_{T}$ to have a simple (e.g., Gaussian) distribution—and then by sampling from the denoising model $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\boldsymbol{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$.
@@
We define the relationship between $\mathbf{y}$ and $\mathbf{z}$ via a forward diffusion process $r(\mathbf{y}|\mathbf{z},\mathbf{x})=r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})=r(\mathbf{y}_{1}|\mathbf{z},\mathbf{x})\prod_{t=1}^{T-1}r(\mathbf{y}_{t+1}|\mathbf{y}_{t},\mathbf{x})$, which transforms $\mathbf{z}$—the latent whose intractable posterior we seek to approximate—into $\mathbf{y}_{T}$, whose posterior is easier to model (possibly conditioned on $\mathbf{x}$). Examples of $r$ include Gaussian forward diffusion processes and discrete noising processes (Austin et al. 2021). The model $q_{\boldsymbol{\phi}}$ is trained to approximately reverse this forward diffusion process.
3.2 Learning Objective: A Markovian ELBO